a marble rolls off the of a table top with an initial speed of 15.0 m/s. it strikes the floor 2.0 from the base of the table. (a) How long ( t = ? ) does it take the marble to reach the floor, what is the height of the table.


1
Expert's answer
2021-07-12T08:42:39-0400
"t=\frac{x}{v}=\frac{2}{15}=0.133\ s\\h=0.5(9.8)(0.133)^2=0.087\ m"